A physics free fall test center includes a 120 m drop tube. This is an evacuated tube which can be used to drop baseball sized objects.
How long is the object in free fall?
ANS:4.95 s (in s)I just need the easiest way to solve it
Answer #1

v^2 - u^2 = 2gs
0 - u^2 = 2*9.8*120

u = 48.497

v = u - gt
0 = 48.497 - 9.8t

or t = 4.948 or 4.95 s